# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ VIII.2 Exponential scaling progress: A mirage or reality?
We have attempted to provide a comprehensive list of all known technical challenges for scaling quantum processors, the goal being to illuminate obstacles that have been overlooked or addressed piecemeal in prior research Alexeev (2024); Bravyi et al. (2022); Humble et al. (2021). Contemporary quantum computing platforms have primarily concentrated on technical hurdles at the hundred-qubit level, constrained by the quality of individual qubits. By anticipating technical scaling challenges from a thousand to a million qubits, our study outlines a holistic system that could provide practical quantum advantage. Not all of these challenges have been addressed in detail in this paper. Our hope is that, by publicizing the obstacles to scaling a quantum supercomputer, we can stimulate the discovery of innovative solutions in both academia and industry. Future revisions of this approach are expected with new discoveries and collaborators.
@@
Acknowledgments. The authors from HPE are supported by the Defense Advanced Research Projects Agency (DARPA) under Air Force Research Laboratory (AFRL) contract no. FA8650-23-3-7313. We would like to thank Eleanor Rieffel for useful discussions and review of the manuscript. Authors from NASA/USRA acknowledge NASA Academic Mission Service contract No. NNA16BD14C. The authors from 1QBit thank our editor, Marko Bucyk, for editorial review of the manuscript. We are grateful to Alexandre Fleury, Einar Gabassov, Mia Kramer, Huy Anh Nguyen, Kevin Nguyen, Katie Olfert, Valentin Senicourt, Yumeng Wang, Chan Woo Yang, and Xiangyi Zhang for useful discussions and support. The authors acknowledge the financial support of Pacific Economic Development Canada (PacifiCan) under project number PC0008525. G. A. M. is grateful for the support of Mitacs. P. R. acknowledges the financial support of Mike and Ophelia Lazaridis, Innovation, Science and Economic Development Canada (ISED), and the Perimeter Institute for Theoretical Physics. Research at the Perimeter Institute is supported in part by the Government of Canada through ISED and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A Analysis of logical circuits
@@
A.1 Workflows for generating logical circuits
In this section, we outline our workflow for generating the logical quantum circuits that serve as input to the resource estimation pipeline, which computes the associated physical resource requirements. The circuits we generate pertain to quantum simulations for estimating the ground-state energy of molecules. We first specify the quantum simulation algorithm used. We then discuss the workflow for how we obtain the quantum circuits that implement this algorithm from the basic specifications of a molecule. Finally, we analyze the various bounds on the errors incurred in the process of generating the logical circuits, and explain how these bounds need to be chosen to ensure the quantum simulations achieve a given target accuracy. While in our study we use the $p$-benzyne molecule as a concrete example, the described methodology applies to other molecules. Our analysis follows closely the approach of Ref. Reiher et al. (2017).
The quantum phase estimation (QPE) algorithm Kitaev (1995b); Abrams and Lloyd (1999) is arguably one of the most rigorous quantum computational approaches for estimating ground-state energies in quantum chemistry. Quantum phase estimation is designed to sample in the eigenbasis of the molecular Hamiltonian $H$ by measuring the phase accumulated on an initial input quantum state acted upon by a unitary operator whose eigenvalue spectrum is a function of the spectrum of $H$. The standard approach is to implement QPE with the time-evolution operator $\exp(-iHt)$. Even more advanced approaches have been proposed; for example, the framework of qubitization Low and Chuang (2019) allows taking a Hamiltonian given by a sum of unitaries (which is the typical case for quantum chemistry Hamiltonians) and constructing a new operation called “qubiterate” that has a functional dependence on the eigenvalues of the Hamiltonian and thus can be used in QPE in place of $\exp(-iHt)$ (see Ref. Berry et al. (2018)). Nevertheless, our quantum resource estimation (QRE) analysis pertains to implementing the standard QPE algorithm, that is, we generate concrete QREs for implementing the time-evolution operator $\exp(-iHt)$ by a quantum circuit. More specifically, for a given molecule, we generate QREs for Hamiltonian simulation based on the use of product formulas (PF). In general, an operator $\mathcal{S}_{p}(t)$ is called an order-$p$ product formula associated with the time-evolution operator $\exp(-iHt)$ for a given Hamiltonian $H$ if (cf. Refs. Childs et al. (2021); Zhuk et al. (2024))
@@ A.1 Workflows for generating logical circuits
Our resource estimation analyses are based on using either the first-order Lie–Trotter formula or the second-order Trotter–Suzuki formula, and the resource estimations pertain to implementing a single Trotter slice based on either of these formulas. In the framework of second quantization, the electronic model Hamiltonian is typically given as
@@
where $\hat{a}^{\dagger}_{p}$ and $\hat{a}_{p}$ are the fermionic creation and annihilation operators, respectively, associated with a given basis set of spin-orbital basis functions $\{\phi_{p}(\bm{x})\}$ (where $\bm{x}\equiv\{\bm{r},\sigma\}$ summarizes the orbital and spin degrees of freedom), and the scalar coefficients $h_{pq}$ and $h_{pqrs}$ are the one- and two-electron integrals, respectively, over the basis functions, computed using the kinetic term and the nuclear and electron–electron coulomb potentials. Numerous software tools exist to derive the second-quantized Hamiltonian from the molecular specifications, which include basic information to fully characterize the system, such as the type of participating atoms and the molecule’s geometry (typically summarized in an $xyz$ file), total charge, and total spin. For this study, we used Tangelo which is an open source Python software package for end-to-end chemistry workflows for quantum computation Senicourt et al. (2022). The $p$-benzyne molecule $\mbox{C}_{6}\mbox{H}_{4}$ (which has zero total charge) exhibits a biradical open-shell singlet ground state (it has zero total spin), with two unpaired electrons. Its geometry is specified by the $xyz$ configuration shown in table 5 (cf.  Section 19 in the supplementary material of Ref. Keller et al. (2015)).
@@
Table 5: Molecular geometry of $p$-benzyne in ångströms, in terms of the $xyz$ file format; cf. Ref. Keller et al. (2015).
@@
In addition to the molecule specifications, we need to select a basis set $\{\phi_{p}(\bm{x})\}$. Basis set selection can be a challenging task. While theoretically an infinite basis is required to represent the true molecular multi-body wavefunction, in practice we cannot perform calculations using an infinite number of basis functions and must therefore rely on using a finite basis set. Numerous basis sets have been introduced and extensively studied in quantum computational chemistry. The most common minimal basis sets are the STO-$n$G basis sets, which are derived from a Slater-type orbital basis set, with $n$ denoting the number of Gaussian primitive functions used to represent each Slater-type orbital. While minimal basis sets are computationally inexpensive, they typically result in insufficiently precise computations. Pople basis sets are a type of split-valence basis sets which use more than one basis function to represent valence orbitals, because it is the valence electrons that typically contribute to the molecular bonding. An entire hierarchy of Pople basis sets have been studied. Importantly, as the basis set grows larger, the resulting approximation gets closer to the true wavefunction; however, increasing the basis set size also results in increasing the required computational resources in space and in time. As a rule of thumb, to achieve semi-quantitative energies, the minimum requirement is to use double-zeta basis sets (such as, e.g., 6-31G or cc-pvdzSenicourt et al. (2022). For our QRE analysis, we used the 6-31G basis set; this basis set yields a good trade-off between accuracy and computation time.
@@
Once a basis set has been selected, we can reduce the size of the system (and thus the computational cost) via active space selection. This concept relies on the notion that, when only considering a subset of the full active space, the resulting loss in correlations affecting the energy computation can be small. For example, in the so-called “frozen-core approximation”, low-lying occupied core orbitals (which typically do not mix with valence orbitals) are “frozen”, that is, they are not included in the computation. Choosing which molecular orbitals to freeze is not a trivial task. Again, we used Tangelo, which provides a means to identify the active space specified by the numbers of molecular orbitals to be included that are energetically next to (i.e., below or above) the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO). For example, the specification “HOMO22-2- 2 and LUMO$+1$” means that we include two additional molecular orbitals below the HOMO and one additional orbital above the LUMO. A common choice is to employ an equal number of additional orbitals to be included next to the HOMO and LUMO; this choice leads to lower energies as opposed to active spaces with unequal numbers of orbitals next to the HOMO and LUMO (cf. Ref. Sennane et al. (2023)). For example, for the $p$-benzyne molecule, the full active space involves 68 active spin-orbitals for the STO-3G basis and 124 active spin-orbitals for the 6-31G, the frozen-core approximation involves 56 active spin-orbitals for the STO-3G basis and 112 active spin-orbitals for the 6-31G, while, for instance, an active space selection ranging from HOMO55-5- 5 to LUMO$+5$ involves only 24 active spin-orbitals for both basis sets. Note that the number of qubits required to encode the system equals the number of active spin-orbitals.
@@
Once both the basis set and the active space have been selected, we can generate the associated second quantized Hamiltonian, as given in eq. 17. The last step is to translate the model Hamiltonian from the second quantization framework to a framework suitable for the quantum circuit model. This step uses a fermion–to–qubit mapping, which is typically either the Jordan–Wigner or the Bravyi–Kitaev transformation, to obtain the Hamiltonian in the Pauli-product form. For a Hamiltonian acting on $n$ qubits, it can be expressed as
@@ A.2 Analysis of Trotter errors and
We now discuss the various errors incurred in the process of generating the logical circuits and how we can guarantee quantum simulations within a given target accuracy by satisfying certain error bounds, which in turn must be within given error budgets. On the logical level, there are three sources of error in implementing the QPE algorithm with Hamiltonian simulation based on using PFs. The first error source is the actual use of an order-$p$ PF, which results in an additive error $\mathcal{O}(t^{p+1})$ in representing the time evolution $\exp(-iHt)$; the additional Trotterization is a technique for dividing the evolution time into many smaller time steps that reduce this error by a constant referred to as “the number of Trotter steps” (or slices). The second error source is associated with circuit synthesis: each term in a PF is implemented by a circuit consisting of single- and two-qubit Clifford gates (such as $H$, $S$, Pauli, and CNOT gates) and some arbitrary-angle, single-qubit rotation $R_{Z}(\theta_{\ell})$ (with angle $\theta_{\ell}$ related to the coefficient $\gamma_{\ell}$ in the Hamiltonian in eq. 18). The latter is approximated by some sequence of the form $HTHST^{\dagger}\dots HS$. This approximation is found using circuit synthesis tools based on either the Solovay–Kitaev (SK) algorithm or the Ross–Selinger (RS) algorithm Selinger (2015); Ross and Selinger (2016) that achieves a more favourable scaling in terms of the $T$ gate count. Either of the circuit synthesis methods incurs an error associated with the approximation. The third error source is associated with the precision of estimating the phase in the actual QPE algorithm. In what follows, we elaborate on these errors and provide useful analytic error bounds, which serve to guarantee that some error budgets in our QRE analysis are satisfied.
@@
We aim to bound $|E_{0}-E_{\mbox{\tiny eff}(t),0}|$, which is the difference between the ground-state energy $E_{0}$ of $H$ (which we aim to estimate) and the lowest eigenvalue of $\tilde{H}_{\mbox{\tiny eff}}(t)$ (which we actually estimate). According to Lemma 3 in the supplementary material of Ref. Reiher et al. (2017), for any given target error bound $\epsilon$, $\|H-H_{\mbox{\tiny eff}}(t)\|\leq\epsilon$ also implies $|E_{0}-E_{\mbox{\tiny eff}(t),0}|\leq\epsilon$. Moreover, according to Lemma 4 in Ref. Reiher et al. (2017), the assumption that $\left\|\exp\left(-itH\right)-\exp\left(-itH_{\mbox{\tiny eff}}(t)\right)\right\|\leq\gamma(t)\,t$ is true for some nondecreasing continuous function $\gamma(t)$ on $[0,\infty)$ implies $\|H-H_{\mbox{\tiny eff}}(t)\|\leq\gamma(t)$. We can use these implications to deduce a relation between the error in energy estimation and the errors associated with the Trotter–Suzuki approximation and the circuit synthesis as follows. Using the triangle inequality multiple times and the analytic bounds given in Equations 21 and 22, we may infer the following:

### v2

@@ IX Concluding remarks
We have outlined the design of a quantum supercomputer that combines the latest advances in HPC and superconducting qubits. A comprehensive list of known technical challenges has been used to illuminate obstacles for scaling quantum processors that have been overlooked or addressed piecemeal in prior research Alexeev (2024); Bravyi et al. (2022); Humble et al. (2021); Svore et al. (2006). In particular, contemporary practical quantum computing investigations have primarily concentrated on technical hurdles at the hundred-qubit level, constrained by the quality of individual qubits. By anticipating technical scaling challenges from hundreds to millions of qubits, our study outlines a holistic system that could provide practical quantum advantage.
For qubit hardware, our thesis is that 300-mm semiconductor processing can radically improve qubit quality, reliability, and scaling. Reproducibility and yield can be greatly improved by replacing junction fabrication from lift-off to a CMOS standard deposition and etch process. Our architecture uses wafer-scale integration, where 20k-qubit modules are connected to 3 Kelvin control circuitry through a 300-mm wiring die. Additionally, an on-chip measurement system eliminates the need for expensive and large volume circulators and pre-amplifiers. Tiling multiple 20k-qubit modules allows fitting 160k qubits in a single DR, and reduces the need for optical interconnects between DRs. This integrated and modular approach thus enables a lower-cost path to scaling.
We introduced a design for a classical supercomputing system that integrates fault-tolerant quantum computing to realize hybrid quantum-classical workloads. HPC is changing, becoming more heterogeneous in response to the need to integrate specialized classical processors. Incorporating quantum processors is just a further step in this trend. We described a hybrid quantum-classical full stack that can be extended as these technologies evolve. The stack is modularized within six layers to allow different software vendors to innovate in parallel. We discussed extensions to existing HPC programming environments and explored various circuit knitting techniques for distributing quantum workloads across multiple QPUs as an alternative to direct quantum networking. We described approaches to designing algorithms that leverage both quantum and classical computing capabilities in close coordination, and discussed extensions to current HPC performance benchmarking standards. At the lower layers of the stack, we detailed FTQC protocols and tight integration of the HPC system with the quantum control digital logic via low-latency and high-bandwidth PCIe-based interfaces. This allows seamless, standardized, and high-performance integration between QPUs, CPUs, GPUs, and any other compute resource in a heterogeneous HPC system. By minimizing communication overheads, our design for a high-performance quantum–classical stack enables calibration, quantum error correction and characterization, adaptive circuit knitting, and any other hybrid process to be performed at the cutting-edge limit of fidelity and runtime.
We also assess our design with detailed resource estimates of classically-hard electronic structure calculations leveraging detailed experimentally-motivated simulation and benchmarking of fault-tolerant quantum computing protocols using superconducting qubits. We showed that for quantum simulations of FeMoco with chemical accuracy ϵ=italic-$\epsilon=$ 1.6 mHa, one requires hundreds of millions of physical qubits and a minimum of 1.4 years to run with current hardware quality (the ‘baseline’ hardware in our study). The runtime can be reduced to approximately four months and the qubit count to tens of millions if the hardware quality is significantly improved to our proposed targets. Our sensitivity analysis revealed that improving the gate fidelities would improve the exponential error suppression of surface codes significantly more than reducing SPAM errors or protecting idling qubits. Another enlightening result of our study is that the logical fidelity of surface code depends considerably on the distribution of qubit errors within the qubit array. Moreover, we demonstrate a threshold behavior at $\sim$6for rows of weaker stabilizer measurements representing capacitive tile-to-tile interconnects, or any anticipated future QPU-to-QPU quantum networking technology (such as optical interconnects). Nevertheless, achieving this fidelity especially at the high Bell-pair generation frequencies required for lattice surgeries between high-distance codes is ambitious for the existing optical interconnect technologies for superconducting qubits. We therefore recommend alternative techniques such as partial fault-tolerant compilation and circuit knitting to be explored in parallel.
Our quantum supercomputing vision ultimately tightly integrates quantum compute, classical compute, and control components at scale using state-of-the-art, yet existing underlying technologies. Not all of the challenges we laid out have been addressed by our approach. Our hope is that, by publicizing the obstacles to scaling quantum computers along with a baseline design, we can stimulate the discovery of innovative solutions in both academia and industry. Future revisions of this approach are expected with new discoveries and collaborators.
@@
Acknowledgments. The authors from HPE are supported by the Defense Advanced Research Projects Agency (DARPA) under Air Force Research Laboratory (AFRL) contract no. FA8650-23-3-7313. PAL and DV were supported in part through the NASA Academic Mission Services (contract NNA16BD14C), as well as the Intelligent Systems Research and Development-3 (ISRDS-3) Contract 80ARC020D0010. The authors thank Marko Bucyk (1QBit) for the editorial review of the manuscript. We are grateful to Alexandre Fleury, Einar Gabassov, Mia Kramer, Huy Anh Nguyen, Kevin Nguyen, Katie Olfert, Valentin Senicourt, Yumeng Wang, Chan Woo Yang, Xiangyi Zhang, Ryan Babbush, Craig Gidney, Matthew Harrigan, N. Cody Jones, Tanuj Khattar, Guang Hao Low, Nicholas Rubin, Akihisa Goban, Masashi Hirose, Shinichi Sunami, Shiro Tamiya, Ming-Zhi Chung, Keisuke Fujii, Yohei Ibe, Toru Kawakubo, Tennin Yan, Andreas Thomasen, Robert Parrish, Bob Sorenson, Jeremy Stevens, Oliver Ezratty, Margaret Martonosi, Taylor Patti, Brucek Khailany, and Michael Brown for useful discussions, valuable feedback, and support. The authors acknowledge the financial support of Pacific Economic Development Canada (PacifiCan) under project number PC0008525. G. A. D. is grateful for the support of Mitacs. P. R. acknowledges the financial support of Mike and Ophelia Lazaridis, Innovation, Science and Economic Development Canada (ISED), and the Perimeter Institute for Theoretical Physics. Research at the Perimeter Institute is supported in part by the Government of Canada through ISED and by the Province of Ontario through the Ministry of Colleges and Universities. Qolab acknowledges financial support from the Development Bank of Japan.
Appendix A Analysis of logic circuits
@@
In this section, we outline our workflow for generating the logical quantum circuits that serve as input to the resource estimation pipeline which computes the associated physical resource requirements. The circuits we generate pertain to quantum simulations for estimating the ground-state energy of molecules. We first specify the quantum simulation algorithm used. We then discuss the workflow for how we obtain the quantum circuits that implement this algorithm from the basic specifications of a molecule. Finally, we analyze the various bounds on the errors incurred in the process of generating the logic circuits, and explain how these bounds must be chosen to ensure that the quantum simulations achieve a given target accuracy. While in our study we use the $p$-benzyne molecule as a concrete example, the described methodology applies to other molecules.
A.1 Workflows for generating logic circuits
The quantum phase estimation (QPE) algorithm Kitaev (1995b); Abrams and Lloyd (1999) is arguably one of the most rigorous quantum computational approaches for estimating ground-state energies in quantum chemistry. Quantum phase estimation is designed to sample in the eigenbasis of the molecular Hamiltonian $H$ by measuring the phase accumulated on an initial input quantum state acted upon by a unitary operator whose eigenvalue spectrum is a function of the spectrum of $H$. The standard approach is to implement QPE with the time-evolution operator $\exp(-iHt)$. More-advanced approaches to electronic-structure quantum simulations are typically based on the framework of qubitization Low and Chuang (2019). This framework allows taking a Hamiltonian given by a sum of unitaries (which is the typical case for quantum chemistry Hamiltonians) and constructing a new operation called “qubiterate” that has a functional dependence on the eigenvalues of the Hamiltonian and thus can be used in QPE in place of $\exp(-iHt)$ (see Ref. Berry et al. (2018)). Our quantum resource estimation (QRE) studies report estimates for two algorithmic approaches: the Trotter-based approach, and the qubitization-based double low-rank factorization algorithm. In what follows, we outline the workflow of QRE for the standard QPE algorithm, with the time-evolution operator $\exp(-iHt)$ implemented using Hamiltonian simulation based on the use of product formulae (PF) and Trotterization. The associated analysis of Trotter errors and their propagation into QPE is discussed in Section A.2. We refer the reader to the literature for various qubitization-based approaches Berry et al. (2019); von Burg et al. (2021a); Lee et al. (2021); Goings et al. (2022); Beverland et al. (2022); Otten et al. (2024a); Bellonzi et al. (2024a); Watts et al. (2024a); Nguyen et al. (2024a); Rocca et al. (2024)). The propagation of errors in qubitization is briefly outlined in Section A.3. A brief description of the double-factorized quantum chemistry simulations used in our QRE studies is given in Appendix C.
In the Trotter-based approaches to QPE, the time-evolution operator $\exp(-iHt)$ is approximated by a PF along with Trotterization. In general, an operator $\mathcal{S}_{p}(t)$ is called an order-$p$ product formula associated with the time-evolution operator $\exp(-iHt)$ for a given Hamiltonian $H$ if (see Refs. Childs et al. (2021); Zhuk et al. (2024))
@@ A.1 Workflows for generating logic circuits
Our resource estimation analyses are based on using either the first-order Lie–Trotter formula or the second-order Trotter–Suzuki formula.
In the framework of second quantization, the electronic model Hamiltonian is typically given as
@@
where $\hat{a}^{\dagger}_{p}$ and $\hat{a}_{p}$ are the fermionic creation and annihilation operators, respectively, associated with a given basis set of spin-orbital basis functions $\{\phi_{p}(\bm{x})\}$ (where $\bm{x}\equiv\{\bm{r},\sigma\}$ summarizes the orbital and spin degrees of freedom), and the scalar coefficients $h_{pq}$ and $h_{pqrs}$ are the one- and two-electron integrals, respectively, over the basis functions, computed using the kinetic term and the nuclear and electron–electron coulomb potentials. Numerous software tools exist to derive the second-quantized Hamiltonian from the molecular specifications, which include the basic information for fully characterizing the system, such as the type of participating atoms and the molecule’s geometry (typically summarized in an $xyz$ file), total charge, and total spin. For this study, we used Tangelo, which is an open source Python software package for end-to-end chemistry workflows for quantum computation Senicourt et al. (2022). The $p$-benzyne molecule $\mbox{C}_{6}\mbox{H}_{4}$ (which has zero total charge) exhibits a biradical open-shell singlet ground state (it has zero total spin), with two unpaired electrons. Its geometry is specified by the xyz file configuration shown in Table 6 (see Section 19 in the supplementary material of Ref. Keller et al. (2015)).
@@
Table 6: Molecular geometry of $p$-benzyne in ångströms, in terms of the xyz file format; see Ref. Keller et al. (2015).
@@
In addition to the molecule specifications, we need to select a basis set $\{\phi_{p}(\bm{x})\}$. Basis set selection can be a challenging task. While theoretically an infinite basis is required to represent the true molecular multi-body wavefunction, in practice we cannot perform calculations using an infinite number of basis functions and must therefore rely on using a finite basis set. Numerous basis sets have been introduced and extensively studied in quantum computational chemistry. The most-common minimal basis sets are the STO-$n$G basis sets, which are derived from a Slater-type orbital basis set, with $n$ denoting the number of Gaussian primitive functions used to represent each Slater-type orbital. While minimal basis sets are computationally inexpensive, they typically result in insufficiently precise computations. Pople basis sets are a type of split-valence basis sets that use more than one basis function to represent valence orbitals, because it is the valence electrons that typically contribute to the molecular bonding. An entire hierarchy of Pople basis sets have been studied. Importantly, as the basis set grows larger, the resulting approximation gets closer to the true wavefunction; however, increasing the basis set size also results in increasing the required computational resources in both space and time. As a rule of thumb, to achieve semi-quantitative energies, the minimum requirement is to use double-zeta basis sets (such as, e.g., 6-31G or cc-pvdzSenicourt et al. (2022). For our QRE analysis, we used the 6-31G basis set; this basis set yields a good trade-off between accuracy and computation time.
@@
Once a basis set has been selected, we can reduce the size of the system (and thus the computational cost) via active space selection. This concept relies on the notion that, when considering only a subset of the full active space, the resulting loss in correlations affecting the energy computation can be small. For example, in the so-called “frozen-core approximation”, low-lying occupied core orbitals (which typically do not mix with valence orbitals) are “frozen”, that is, they are not included in the computation. Choosing which molecular orbitals to freeze is not a trivial task. Again, we used Tangelo, which provides a means to identify the active space specified by the numbers of molecular orbitals to be included that are energetically next to (i.e., below or above) the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO). For example, the specification “HOMO22-2- 2 and LUMO$+1$” means that we include two additional molecular orbitals below the HOMO and one additional orbital above the LUMO. A common choice is to employ an equal number of additional orbitals to be included next to the HOMO and LUMO; this choice leads to lower energies as opposed to active spaces with unequal numbers of orbitals next to the HOMO and LUMO (see Ref. Sennane et al. (2023)). For example, for the $p$-benzyne molecule, the full active space involves 68 active spin-orbitals for the STO-3G basis and 124 active spin-orbitals for the 6-31G basis, and the frozen-core approximation involves 56 active spin-orbitals for the STO-3G basis and 112 active spin-orbitals for the 6-31G basis, while, for instance, an active space selection ranging from HOMO55-5- 5 to LUMO$+5$ involves only 24 active spin-orbitals for both basis sets. Note that the number of qubits required to encode the system equals the number of active spin-orbitals.
@@
Once the basis set and the active space have been selected, we can generate the associated second quantized Hamiltonian, as given in Equation 19. The last step is to translate the model Hamiltonian from the second quantization framework to a framework suitable for the quantum circuit model. This step uses a fermion–to–qubit mapping, which is typically either the Jordan–Wigner Jordan and Wigner (1993) or the Bravyi–Kitaev Bravyi and Kitaev (2002) transformation, to obtain the Hamiltonian in the Pauli-product form. For a Hamiltonian acting on $n$ qubits, it can be expressed as
@@ A.2 Analysis of Trotter errors and
We now discuss the various errors incurred in the process of generating the logic circuits in the Trotter-based approaches to QPE and how we can guarantee electronic-structure quantum computations to estimate the Hamiltonian eigenvalues within a given target accuracy by satisfying certain error bounds, which in turn must be within given error budgets. Our analysis of Trotter errors and their propagation into phase estimation closely follows the approach of Ref. Reiher et al. (2017). The propagation of errors in qubitization is discussed in the next subsection.
On the logical level, there are three sources of error in implementing the QPE algorithm with Hamiltonian simulation based on using PFs. The first error source is the actual use of an order-$p$ PF, which results in an additive error $\mathcal{O}(t^{p+1})$ in representing the time evolution $\exp(-iHt)$; the additional Trotterization is a technique for dividing the evolution time into many smaller time steps that reduce this error by a constant referred to as “the number of Trotter steps” (or slices). The second error source is associated with circuit synthesis: each term in a PF is implemented by a circuit consisting of single- and two-qubit Clifford gates (such as $H$, $S$, Pauli, and CNOT gates) and some arbitrary-angle, single-qubit rotation $R_{Z}(\theta_{\ell})$ (with angle $\theta_{\ell}$ related to the coefficient $\gamma_{\ell}$ in the Hamiltonian in Equation 20). The latter is approximated by some sequence of the form $HTHST^{\dagger}\dots HS$. This approximation is found using circuit synthesis tools based on either the Solovay–Kitaev (SK) algorithm or the Ross–Selinger (RS) algorithm Selinger (2015); Ross and Selinger (2016) that achieves a more favourable scaling in terms of the $T$ gate count. Either of the circuit synthesis methods incurs an error associated with the approximation. The third error source is associated with the precision of estimating the phase in the actual QPE algorithm. In what follows, we elaborate on these errors and provide useful analytic error bounds, which serve to guarantee that some error budgets in our QRE analysis are satisfied.
@@
We aim to bound $|E_{0}-E_{\mbox{\tiny eff}(t),0}|$, which is the difference between the ground-state energy $E_{0}$ of $H$ (which we aim to estimate) and the lowest eigenvalue of $\tilde{H}_{\mbox{\tiny eff}}(t)$ (which we actually estimate). According to Lemma 3 in the supplementary material of Ref. Reiher et al. (2017), for any given target error bound $\epsilon$, $\|H-H_{\mbox{\tiny eff}}(t)\|\leq\epsilon$ also implies $|E_{0}-E_{\mbox{\tiny eff}(t),0}|\leq\epsilon$. Moreover, according to Lemma 4 in Ref. Reiher et al. (2017), the assumption that $\left\|\exp\left(-itH\right)-\exp\left(-itH_{\mbox{\tiny eff}}(t)\right)\right\|\leq\gamma(t)\,t$ is true for some nondecreasing continuous function $\gamma(t)$ on $[0,\infty)$ implies $\|H-H_{\mbox{\tiny eff}}(t)\|\leq\gamma(t)$. We can use these implications to deduce a relation between the error in energy estimation and the errors associated with the Trotter–Suzuki approximation and the circuit synthesis as follows. Using the triangle inequality multiple times and the analytic bounds given in Equations 23 and 24, we may infer the following: